hi can you help me solve this twp step equation: t/7 -5 =-10
Answer by checkley79(3341)

t/7 -5 =-10 ADD 5.
t/7=-10+5
t/7=-5 CROSS MULTIPLY.
t=7*-5
t=-35 ans.
Proof:
-35/7-5=-10
-5-5=-10
-10=-10